Circular No. 8655

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005my, 2005mz, 2005na Three apparent supernova have been found on unfiltered CCD frames:

2005my by P. Luckas, Perth, W. Australia; O. Trondal, Oslo, Norway; and M. Schwartz, Patagonia, AZ (0.35-m Tenagra telescope at Perth; cf. IAUC 8616); and 2005mz and 2005na by J. Newton and T. Puckett (both with a 0.35-m reflector telescope at Portal, AZ; 2005mz with M. Peoples and 2005na with D. Ceravolo; cf. IAUC 8651).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005my Dec. 30.55 4^{h}01^{m}53.^{s}13 −41^{o}56′08′′.3 17.8 15′′.9 E, 23′′.0 N
2005mz Dec. 31.25 3 19 49.88 +41 30 18.6 18.2 19′′.2 E, 23′′.6 S
2005na Dec. 31.35 7 01 36.62 +14 07 59.7 16.1 1′′.6 W, 7′′.4 S

Additional magnitudes for 2005my in ESO 302-G27: 2005 Oct. 14.75 UT, [18.5 (Luckas); 2006 Jan. 1.53, 17.3. Additional magnitudes for 2005mz in NGC 1275: 2005 Dec. 3, [19.8 (Puckett); 2006 Jan. 1.13, 18.0 (Puckett, 0.60-m reflector). Additional magnitudes for 2005na in UGC 3634: 2005 Nov. 15, [19.7 (Puckett); 2006 Jan. 2.24, 15.6 (Puckett, 0.35-m reflector). SNe 2005mz and 2005na are type-Ia supernovae (cf. CBET 351).

V838 MONOCEROTIS A. Henden, AAVSO; and U. Munari and H. Navasardyan, Astronomi-

cal Observatory, Padova, write: “Our photometric and optical spectroscopic monitoring shows V838 Mon to be smoothly continuing along its slow evolution path. UBV (RI)_{c} CCD photometry obtained on 2005 Dec. 25.33 UT with the U.S. Naval Observatory 1.0-m telescope at Flagstaff Station gives V = 15.44 (+0.03), U–B = +0.07 (+0.02), B–V = +1.17 (−0.03), V–R_{c} = +1.96 (−0.18), R_{c}–I_{c} = +2.71 (+0.02), with quantities in parentheses being the difference compared with values obtained with the same instrumentation one year ago. Near-infrared photometry secured with the Flagstaff Station 1.55-m telescope on Dec. 18.3 gives J = 7.51, H = 6.45, K′ = 5.70, continuing a steady, slow decline that has been in progress since November 2003. Spectra (range 420–780 nm; dispersion 0.4 nm) obtained with the Asiago 1.82-m telescope (+ AFOSC) on various dates in December 2005 closely resemble those of the last two years, dominated by extremely strong VO and TiO absorption bands of the L-type supergiant in the red and the continuum of the B3V companion in the blue. This seems to be in contradiction to the dramatic fading and spectral changes mentioned by Lynch et al. on IAUC 8645.”

2006 January 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green